# Formatting Instructions for ICLR 2025 Conference Submissions

### v1

@@ B.1 Mutation Rate (Depth Pruning)
The mutation rate plays a crucial role in balancing exploration and exploitation. A higher mutation rate allows for broader exploration of the search space; however, this space grows exponentially with the number of mutations. As a result, when trying to approach the optimum in terms of Hamming distance, the proportion of “good” offspring decreases significantly with an increasing mutation rate. Consequently, in a smooth fitness landscape, we expect significantly faster optimization with a lower mutation rate. This intuition for the mutation rate is fundamentally different from that of a learning rate in continuous optimization.
@@
A widely used strategy for balancing the advantages and disadvantages of different mutation rates involves self-adjusting mutation rates, which have been shown to be effective both theoretically and in practice (Kern et al., 2004; Doerr et al., 2019). These methods decrease the mutation rate when progress is relatively ”easy”, and increase it when progress becomes difficult, offering a greater chance of escaping local optima.
@@ B.2 Multi-Step Selection (Unstructured Sparsity)
In a similar vein to the discussion in Appendix B.1, the number of offspring can also be dynamically adapted. Ideally, the number of offspring should increase to the point where the computational effort is compensated by the number of generations, as outlined in Theorem 1. Methods such as the Self-Adjusting $(1,\lambda)$-EA have recently gained significant theoretical interest and have been shown to automatically determine “ideal” offspring sizes on specific problems (Hevia Fajardo & Sudholt, 2021; Kaufmann et al., 2022). Although we have not experimented with such adaptive methods, we see significant potential for future work in this area, particularly considering the multi-step selection we employ.
@@ B.3 Fitness Environment (Quantization)
Appendix C Hyperparameter Setting
Here, we provide an overview of the hyperparameters used in our experiments. As shown in Table 8, different hyperparameters were employed for different applications due to the varying nature of their search spaces. Across all applications, we sampled the number of mutations from the distribution $\min(U_{1},U_{2})$ with $U_{1},U_{2}\sim\text{Uniform}(1,3)$, which closely mimics the behavior of using only one mutation (see the ablation study in Appendix B.1).
For Depth Pruning, where each block has only two choices and significantly fewer blocks are present compared to layers in other methods, we leveraged the insight from Theorem 1, which suggests that the number of required generations scales proportionally to $k(n-k)$, where $k$ represents the number of removed blocks and $n$ the total number of blocks.
For Unstructured Sparsity, the search space is considerably larger, with more than 10 choices per layer [^5] If necessary, the step size can be increased and the number of compression levels reduced with minimal impact on performance. We opted for a higher number of compression levels for a fair comparison with OWL.. As a result, more generations are necessary to converge because each generation only makes small improvements in terms of Hamming distance from the optimum.
@@ Appendix C Hyperparameter Setting
For all these applications, we adopted a conservative approach for the number of generations to better understand convergence. In practice, we need significantly fewer generations to converge, as displayed in Figure 3, Figure 5 and Figure 10. Additionally, we showed a “super-fast” version (in terms of time per iteration) that is significantly more lightweight.
@@ D.1 Full Results
Table 9: Depth pruning of Mistral-7B-v0.3.
[…]
0% Dense 4.82 7.72 6.41
12.5% EvoPress 6.06 9.00 7.42
EvoPress (Attn.+MLP) 6.33 9.44 7.80
ShortGPT 7.19 10.18 8.46
Cosine Similarity (Window) 7.19 10.18 8.46
Weight Subcloning 7.19 10.18 8.46
Shortened Llama 6.64 9.71 7.94
25% EvoPress 8.66 12.04 9.92
EvoPress (Attn.+MLP) 9.46 13.02 10.59
ShortGPT 43.26 40.16 29.54
Cosine Similarity (Window) 33.75 54.07 36.26
Weight Subcloning 43.26 40.16 29.54
Shortened Llama 14.94 19.30 14.73
37.5% EvoPress 17.52 21.60 16.90
EvoPress (Attn.+MLP) 21.62 25.17 18.97
ShortGPT 2898.98 2722.66 981.99
Cosine Similarity (Window) 1034.09 2471.86 1050.56
Weight Subcloning 2898.98 2722.66 981.99
Shortened Llama 440.20 442.09 486.15
50% EvoPress 61.75 54.15 43.23
EvoPress (Attn.+MLP) 108.91 99.74 69.07
ShortGPT 2422.72 2134.92 1083.51
Cosine Similarity (Window) 3411.47 1934.16 1740.91
Weight Subcloning 2422.72 2134.92 1083.51
Shortened Llama 5241.76 3595.71 1953.14
@@
Here, we present our additional results for depth pruning experiments on Mistral-7B-v0.3 (Table 9), Llama-2-7B (Table 10), Llama-3-8B (Table 11), and Llama-3.1-8B (Table 12). Across all levels of sparsities, EvoPress consistently outperforms previous methods. Additionally, Table 9 includes results where only entire transformer blocks are removed by EvoPress, showcasing that the significant gains are not primarily due to this relaxation, and that our method performs better than baselines even when dealing with this coarser search space.
@@ D.2 Practical Convergence
EvoPress not only produces superior compression configurations, it also does so extremely efficiently. Figure 5 displays that the evolutionary search produces better compressed models than previous techniques in a matter of minutes, with full convergence in around half an hour.
@@ D.3 Locality of Dropped Blocks
Table 13: First 16 blocks in removal order of ShortGPT, Weight Subcloning and Shortened Llama on three different models.
@@ D.4 Correlation of Scores with Perplexity
In this experiment, we first calculated the cosine similarity and squared error for each block by comparing activations before and after the block. Next, we randomly removed subsets of blocks (excluding the first and last two) and for each configuration, computed the average cosine similarity and squared error. The results are shown in Figure 7. Initially, the average squared error exhibited a negative correlation, as the $\ell_{2}$-norm of the activations increased with depth. This led to configurations with early blocks removed having small average error. To mitigate this, we normalized the activations prior to computing the squared error, which significantly improved the correlation, resulting in performance comparable to cosine similarity. However, as sparsity increased, the correlation degraded significantly for both methods, offering insight into why removal techniques based on scoring fail even at moderate levels of sparsity. The experiment was done using 131,072 tokens from the Fineweb-Edu calibration dataset.
Figure 7: Effect of removing random subsets of blocks for Llama-3-8B.
@@ Appendix E Additional Unstructured Sparsity Results
Table 14: Performance of various sparsity profiles at 50% sparsity
@@
Table 15: Performance of various sparsity profiles at 60% sparsity
@@ Appendix F Additional Quantization Results
Table 16: Performance of EvoPress on 2.25 bit and 2.5 bit quantization
@@ F.2 Practical Convergence
Similar to unstructured sparsity, EvoPress also demonstrates rapid convergence when applied to quantization. As shown in Figure 10, the majority of improvements occur within two GPU hours, with full convergence achieved after approximately eight GPU hours. If needed, this optimization time could be further shortened by tuning the hyperparameters, similarly to the “super-fast” version for unstructured sparsity discussed in Section 4.2. However, we observed that the convergence dynamics are less smooth compared to unstructured sparsity, likely due to the limited number of quantization levels available (practically only 2, 3, and 4 bit are used), resulting in a less smooth fitness landscape.

### v2

@@ B.1 Mutation Rate (Depth Pruning)
The mutation rate plays a crucial role in balancing exploration and exploitation. A higher mutation rate allows for broader exploration of the search space; however, this space grows exponentially with the number of mutations. As a result, when trying to approach the optimum in terms of Hamming distance, the proportion of “good” offspring decreases significantly with an increasing mutation rate. Consequently, in a smooth fitness landscape, we expect significantly faster optimization with a lower mutation rate.
@@
A widely used strategy for balancing the advantages and disadvantages of different mutation rates involves self-adjusting mutation rates, which have been shown to be effective both theoretically and in practice (Kern et al., 2004; Doerr et al., 2019). These methods decrease the mutation rate when progress is relatively “easy”, and increase it when progress becomes difficult, offering a greater chance of escaping local optima.
@@ B.2 Multi-Step Selection (Unstructured Sparsity)
In a similar vein to the discussion in Appendix B.1, the number of offspring can also be dynamically adapted. Ideally, the number of offspring should increase to the point where the computational effort is compensated by the number of generations, as outlined in Theorem 3.1. Methods such as the Self-Adjusting $(1,\lambda)$-EA have recently gained significant theoretical interest and have been shown to automatically determine “ideal” offspring sizes on specific problems (Hevia Fajardo & Sudholt, 2021; Kaufmann et al., 2022).
@@ B.3 Fitness Environment (Quantization)
Appendix C Experimental Setup
C.1 Hyperparameter Setting
Here, we provide an overview of the hyperparameters used in our experiments. As shown in Table 8, we employed different choices for the number of tokens, offspring, and generations for different applications to account for the size of the respective search space. However, the search is very robust with respect to these choices, and using one set of hyperparameters for another application yields similar results. This is also demonstrated through the “super-fast” version, which uses drastically different hyperparameters, but achieves comparable performance (see Figure 3 in main text).
Across all applications, we sampled the number of mutations from the distributions $\min(U_{1},U_{2})$ with $U_{1},U_{2}\sim Unif(1,3)$, which closely mimics the behavior of using only one mutation (see the ablation study in Appendix 4).
For Depth Pruning, where each block has only two choices and significantly fewer blocks are present compared to layers in other methods, we leveraged the insight from Theorem 3.1, which suggests that the number of required generations scales proportionally to $k(n-k)$, where $k$ represents the number of removed blocks and $n$ the total number of blocks.
For Unstructured Sparsity, the search space is considerably larger, with more than 10 choices per layer [^4] If needed, one could increase the step size and reduce the number of compression levels to load.. As a result, more generations are necessary to converge because each generation only makes small improvement in terms of Hamming distance from the optimum.
@@ C.1 Hyperparameter Setting
For all these applications, we adopted a conservative approach to the number of generations to better understand convergence. In practice, we need significantly fewer generations to converge close to optimum, as demonstrated in Section 4.2, Appendix D.2, Appendix F.2, and Appendix B.3. Additionally, we showed a much faster version (in terms of time per iteration) that uses significantly less tokens and has fewer offspring.
@@
C.2 Robustness to Random Seed
To evaluate the robustness of EvoPress, we conducted 16 independent runs with different random seeds. Specifically, we used the “super-fast” variant to determine the optimal compression allocation for Llama-3-8B at 70% sparsity, assessing perplexity scores on the C4, Wikitext2, and hold-out Fineweb-Edu datasets. The results indicate that EvoPress is highly robust, as reflected by the low standard deviation observed across the hold-out metrics (Figure 7). For example, after 1000 generations of the “super-fast” variant, the configurations found achieve a mean C4 perplexity of 33.82 with a standard deviation of 0.61, compared to 52.32 for the next best method, OWL. This improvements achieved by EvoPress are therefore statistically relevant. Furthermore, as shown in Figure 6, the configurations identified across different runs are highly similar, which is expected to improve further with additional generations.
Figure 6: Configurations identified by EvoPress on Llama-3-8B after 1000 generations show high similarity across different seeds. The y-axis represents the depth of the respective transformer block, while the x-axis denotes the corresponding layer (q: query, k: key, v: value, o: output, u: MLP up, g: MLP gate, d: MLP down).
Figure 7: Convergence behavior of the “super-fast” variant across 16 independent runs. The extremely low standard deviation (shaded area) demonstrates the robustness of the method, which suggests that local optima do not pose significant challenges to the search.
@@ D.1 Full Results
Here, we present our additional results for depth pruning experiments on Mistral-7B-v0.3 (Table 12), Llama-2-7B (Table 9), Llama-3-8B (Table 10), and Llama-3.1-8B (Table 11). Across all levels of sparsities, EvoPress consistently outperforms previous methods. Additionally, Table 12 includes results where only entire transformer blocks are removed by EvoPress. This showcases that the significant gains are not primarily due to this relaxation, and that our method performs better than baselines even when dealing with this coarser search space.
@@
Table 12: Depth pruning of Mistral-7B-v0.3.
[…]
0% Dense 4.82 7.72 6.41
12.5% EvoPress 6.06 9.00 7.42
EvoPress (Attn.+MLP) 6.33 9.44 7.80
ShortGPT 7.19 10.18 8.46
Cosine Similarity (Window) 7.19 10.18 8.46
Weight Subcloning 7.19 10.18 8.46
Shortened Llama 6.64 9.71 7.94
25% EvoPress 8.66 12.04 9.92
EvoPress (Attn.+MLP) 9.46 13.02 10.59
ShortGPT 43.26 40.16 29.54
Cosine Similarity (Window) 33.75 54.07 36.26
Weight Subcloning 43.26 40.16 29.54
Shortened Llama 14.94 19.30 14.73
37.5% EvoPress 17.52 21.60 16.90
EvoPress (Attn.+MLP) 21.62 25.17 18.97
ShortGPT 2898.98 2722.66 981.99
Cosine Similarity (Window) 1034.09 2471.86 1050.56
Weight Subcloning 2898.98 2722.66 981.99
Shortened Llama 440.20 442.09 486.15
50% EvoPress 61.75 54.15 43.23
EvoPress (Attn.+MLP) 108.91 99.74 69.07
ShortGPT 2422.72 2134.92 1083.51
Cosine Similarity (Window) 3411.47 1934.16 1740.91
Weight Subcloning 2422.72 2134.92 1083.51
Shortened Llama 5241.76 3595.71 1953.14
@@ D.2 Practical Convergence
EvoPress identifies superior compression profiles in a highly efficient manner. Figure 8 displays that the evolutionary search produces better compressed models than previous techniques in a matter of minutes, with full convergence in around half an hour.
@@ D.3 Locality of Dropped Blocks
Table 13: First 16 blocks in removal order of ShortGPT, Weight Subcloning and Shortened Llama on different models.
@@ D.4 Correlation of Scores with Perplexity
In this experiment, we first calculated the cosine similarity and squared error for each block by comparing activations before and after the block. Next, we randomly removed subsets of blocks (excluding the first and last two) and for each configuration, computed the average cosine similarity and squared error. The results are shown in Figure 11. Initially, the average squared error exhibited a negative correlation, as the $\ell_{2}$-norm of the activations increased with depth. This led to configurations with early blocks removed having small average error. To mitigate this, we normalized the activations prior to computing the squared error, which significantly improved the correlation, resulting in performance comparable to cosine similarity. However, as sparsity increased, the correlation degraded significantly for both methods, offering insight into why removal techniques based on scoring fail even at moderate levels of sparsity. Meanwhile, when removing only a small number of blocks, the average perplexity when removing each block separatly is a strong predictor of the performance after removing an entire set of blocks, as depicted in Figure 10. We conclude that error monotonicity holds at smaller compression levels, but decays rapidly at medium sparsity. The experiments were done using 131,072 tokens from the Fineweb-Edu calibration dataset.
Figure 10: Effect of removing random subsets of $2$, $4$, and $6$ blocks for Llama-3-8B. The x-axis depicts the average perplexity when removing each block separately, while the y-axis depicts the perplexity after removing the entire set of blocks. Error monotonicity holds for smaller compression levels, but decays with increasing sparsity.
Figure 11: Effect of removing random subsets of $8$, $10$, and $12$ blocks for Llama-3-8B. The x-axis depicts the average cosine similarity (first row), average squared error (second row), and the average squared error after normalization (third row) when removing each block separately, while the y-axis depicts the perplexity after removing the entire set of blocks. Error monotonicity breaks at medium sparsity rates, providing an explanation for the failure of scoring-based methods at these sparsity levels.
@@ Appendix E Additional Unstructured Sparsity Results
Table 14: Performance of various sparsity profiles at 50% sparsity.
@@
Table 15: Performance of various sparsity profiles at 60% sparsity.
@@ Appendix F Additional Quantization Results
Table 16: Performance of EvoPress on 2.25 bit and 2.5 bit quantization.
@@ F.2 Practical Convergence
Similar to unstructured sparsity, EvoPress also demonstrates rapid convergence when applied to quantization. As shown in Figure 14, the majority of improvements occur within two GPU hours, with full convergence achieved after approximately eight GPU hours. If needed, this optimization time could be further shortened by tuning the hyperparameters, similarly to the “super-fast” version for unstructured sparsity discussed in Section 4.2. However, we observed that the convergence dynamics are less smooth compared to unstructured sparsity, likely due to the limited number of quantization levels available (practically only 2, 3, and 4 bit are used), which results in a less smooth fitness landscape.
@@ F.3 Quantization Profiles
Appendix G Running Time Comparisons
In our main results, we did not include a stopping criteria into the optimization process, but instead continued the search well beyond convergence to observe the convergence behavior. Here, we include a simple stopping criteria to our search, which terminates the search whenever the perplexity on the training set has not improved in the last 20 generations. Based on this version we compare the runtime requirements of EvoPress with baseline methods. However, it should be noted that the amount of data used impacts both runtime and compression quality (with specific trade-offs being method-dependent), meaning one can generally trade-off both objectives, which makes runtime comparison difficult.
In Table 17 we compare EvoPress for block dropping to various baseline methods on Llama-2-7B, where we, additionally to the main results, also compare against the iterative search methods BlockPruner (Zhong et al., 2024) and SLEB (Song et al., 2024). As the results show, EvoPress achieves the best compressed models across all sparsity ratios, while being quicker than other search methods. Scoring-based methods are considerably faster than the search-based methods, but produce significantly worse models.
Table 17: Comparison of various depth pruning methods at different sparsities. We report the runtime on a single NVIDIA RTX 4090 GPU. Note that some methods perform forward passes on a strongly depth-pruned model, which explains the runtime difference for EvoPress on 50% and 12.5% sparsity.
[…]
0% - 6.40 5.12 6.99 - -
12.5% ShortGPT 9.30 8.86 10.78 46s 65K
12.5% Cosine Similarity (Window) 8.51 7.53 9.82 46s 65K
12.5% Weight Subcloning 9.60 9.09 11.06 49s 65K
12.5% ShortenedLlama 8.57 7.68 10.44 8min 2.1M
12.5% SLEB 7.71 6.51 8.78 70min 18.8M
11.4% BlockPruner 7.88 7.88 8.86 100min 26.6M
12.5% EvoPress 7.61 6.41 8.67 58min 12.4M
25% ShortGPT 21.16 23.41 30.30 44s 65K
25% Cosine Similarity (Window) 17.37 16.60 21.04 46s 65K
25% Weight Subcloning 21.16 23.41 30.30 49s 65K
25% ShortenedLlama 11.81 13.86 14.08 8min 2.1M
25% SLEB 10.21 9.45 12.03 122min 35.2M
25% BlockPruner 11.31 12.49 13.32 152min 42.8M
25% EvoPress 10.06 10.31 11.99 44min 10.0M
37.5% ShortGPT 54.07 70.94 63.51 44s 65K
37.5% Cosine Similarity (Window) 151.10 192.07 212.60 45s 65K
37.5% Weight Subcloning 54.07 70.94 63.51 49s 65K
37.5% ShortenedLlama 20.37 35.37 26.07 8min 2.1M
37.5% SLEB 17.29 19.57 20.91 159min 49.1M
36.4% BlockPruner 19.41 28.75 22.44 185min 54.8M
37.5% EvoPress 16.09 17.44 19.87 53min 19.7M
50% ShortGPT 180.51 226.14 171.04 44s 65K
50% Cosine Similarity (Window) 3611.06 4570.15 2876.83 45s 65K
50% Weight Subcloning 180.51 226.14 141.04 49s 65K
50% ShortenedLlama 68.79 145.78 87.40 8min 2.1M
50% SLEB 51.30 117.21 58.56 184min 60.1M
48.9% BlockPruner 52.28 97.23 60.47 208min 64.9M
50% EvoPress 34.20 47.15 40.10 31min 12.4M
In Table 18 we provide a comparison of EvoPress with BESA (Xu et al., 2024), a method for non-uniform sparsity allocation. BESA performs gradient-descent based search for rowwise sparsity allocation using straight-through-estimation (Bengio et al., 2013), with uniform sparsity on a per-block level. Therefore, both methods are not fully comparable, as EvoPress searches for layerwise sparsities globally (across the entire model, not just blockwise), while keeping per-row sparsities uniform. We demonstrate that both methods are complementary by performing a two-stage search, first, for blockwise sparsities using EvoPress, then, for rowwise sparsities with fixed per-block sparsity using BESA. This merged version produces best results on 50% and 60% sparsity, while it performs slightly worse than pure EvoPress on 70% sparsity. For a fair comparison, we search on top of a Wanda (Sun et al., 2024) layer database for all methods. The runtime requirements for both BESA and EvoPress are similar, as there is an additional hyperparameter sweep for BESA, which is not provided in the author’s implementation (authors provide tuned coefficients). Similarly, we have not included OWL in this comparison, as the authors have not provided instructions on how hyperparameters were obtained.
Table 18: Comparison of EvoPress with uniform pruning and BESA on Llama-2-7B. Both EvoPress as well as BESA (layerwise) operate on a layer database generated via Wanda. BESA (rowwise) additionally learns per-output-channel sparsities, which target a different dimension of the compression space than EvoPress’s layerwise allocation. EvoPress+BESA applies BESA using the average per-block sparsities found from running EvoPress, showing that both methods can work complementary. Here, we use the lightweight version of EvoPress with early stopping.
[…]
50.00 Uniform N/A 7.60 6.41 8.97 -
49.79 BESA Layerwise 7.55 6.32 8.93 16min*
50.00 EvoPress Layerwise 7.48 6.28 8.79 87min
49.70 BESA Rowwise 7.39 6.19 8.72 33min*
49.72 EvoPress+BESA Rowwise 7.37 6.16 8.71 120min
60.00 Uniform N/A 10.22 9.56 12.95 -
59.46 BESA Layerwise 9.82 8.80 12.38 16min*
58.91 EvoPress Layerwise 8.75 8.03 11.94 102min
59.99 EvoPress Layerwise 9.01 8.66 10.96 94min
58.90 BESA Rowwise 9.15 8.11 11.88 33min*
59.12 EvoPress+BESA Rowwise 8.59 7.58 10.58 135min*
70.00 Uniform N/A 64.02 136.53 97.53 -
69.97 BESA Layerwise 52.65 66.51 67.58 16min*
68.52 EvoPress Layerwise 13.53 15.61 16.90 111min
69.99 EvoPress Layerwise 15.55 24.04 19.54 120min
68.51 BESA Rowwise 16.53 17.79 24.09 33min*
68.03 EvoPress+BESA Rowwise 13.83 14.19 19.52 144min*
  • Sparsity values in red are more than 0.5% below the target. They are a result of BESA only enforcing a soft constraint on the effective sparsity.
    BESA runtimes do not include the hyperparameter sweep over ‘-d-coef‘.